# Recent Advances in Discrete Speech Tokens: A Review

### v1

@@ I Introduction
The rapid advancement of large language models (LLMs) in natural language processing has revolutionized speech generation tasks [1, 2], with speech being tokenized and modeled using decoder-only Transformers [3]. Recent efforts, such as GSLM [4] and AudioLM [5], aim to develop text-free spoken LLMs, akin to how current LLM-powered chatbots enable text-based interactions. Other works, including VALL-E [6] and VioLA [7], extend this approach to downstream conditional speech generation tasks, such as zero-shot text-to-speech generation and speech translation. However, this paradigm requires data to be tokenized, as typical LLMs can only process discrete data. Textual tokens naturally meet this requirement because they are designed as discrete units separated by clear boundaries, whereas raw speech signals are continuous and boundary-less. Therefore, a necessary step before applying speech data to LLM is the tokenization of speech, which aims to:
To convert long speech waveforms into short discrete tokens for downstream tasks. These tokens should be compatible with the underlying textual representations, especially with language modeling approaches targeted at speech.
@@
As a result, significant efforts have been directed towards developing more efficient and powerful speech tokenization methods. Generally, these methods are based on two distinct principles, giving rise to two types of speech tokens: acoustic tokens and semantic tokens. Acoustic tokens are derived from neural codecs designed to encode speech at a low bitrate while preserving as much information as possible. In contrast, semantic tokens originate from speech self-supervised learning (SSL) [8], which aims to transform the speech representation into a more phonetic or semantic space, making it easier for speech recognition. These two nearly independent lines of research magically intersect in the context of language modeling for speech. Now, there are also efforts that try to design a speech tokenizer that accomplishes the two objectives simultaneously [9, 10]. Consequently, speech tokenization has become a core problem of speech processing under the new paradigm, especially for speech generation and spoken language modeling.
Despite the rapid development and numerous recent works, a comprehensive taxonomy of methodologies in discrete speech tokens has not been clearly constructed. Existing reviews [8, 11, 12, 1, 13, 2] in this field often overlook the diverse categories and methodologies in both acoustic and semantic tokens. For example, [1, 2] focus primarily on methods in spoken language modeling, providing only brief descriptions of some speech tokens used in existing models. Wu et al. [12] introduces several neural audio codecs but largely ignore the realm of semantic tokens. In this review, we provide a comprehensive overview of the concepts, methods, and characteristics of various types of discrete speech tokens, with their applications in spoken language understanding, speech generation, and spoken dialogue models. We hope that through this review, the community can have a clear understanding of the current development and key technologies of discrete speech tokens, so as to promote further research in the future.
@@
  • This review is the first to focus specifically on discrete speech tokens with sufficient depth in the era of speech large language models (LLMs).
@@
The structure of this review is shown in Fig.2. We will first briefly introduce the technical pre-requisites of discrete speech tokens in Section II. Following [5, 14, 15], we classify discrete speech tokens into acoustic and semantic tokens based on their purposes as two major types. We will characterize the two major types of tokens in Section III and IV respectively. In Section V, we summarize the existing benchmarks and analyses of discrete speech tokens, and conduct a comparison of all kinds of tokens on their reconstruction and voice conversion performance. This helps understand the different characteristics of each kind of tokens. Section VI briefly presents several downstream application paradigms of discrete speech tokens besides compression and transmission. Section VII concludes the current challenges and outlook for the future development of discrete speech tokens.
@@ II Pre-requisites: Discrete Representation Learning
Discrete speech tokens are obtained through the quantization of continuous speech representations. In the deep learning era, this quantization is usually achieved by offline clustering or online vector quantization algorithms. This section provides a concise overview of these established quantization methodologies commonly used in discrete speech tokens.
Denote $\bm{x}\in\mathbb{R}^{d}$ as a vector in the $d$-dimensional continuous space. A quantization process $q$ transforms $\bm{x}$ into a discrete token in a finite set, i.e. $q(\bm{x}):\mathbb{R}^{d}\to\{1,2,...,V\}$ where $V$ is the vocabulary size. The output tokens are sometimes referred to as indexes in the finite $V$-cardinal set. The function $q$ is usually associated with a codebook $\mathcal{C}=\{\bm{c}_{1},\bm{c}_{2},...,\bm{c}_{V}\}$ where every code-vector $\bm{c}_{i}\in\mathbb{R}^{d}$ corresponds to the $i$-th token. The code-vectors are representations of tokens in the original $d$-dimensional space. As $V$ integers can be encoded using $\lceil\log_{2}V\rceil$ raw bits [^1] We use $\lceil z\rceil$ to denote the ceiling of a scalar $z$, i.e., the smallest integer greater than or equal to $z$. Similarly, $\lfloor z\rfloor$ denotes the floor of $z$, i.e., the largest integer less than or equal to $z$., quantization often compresses the cost for data storage and transmission to a great extent.
@@ II-A Offline Clustering
Clustering is a simple approach for quantization. Given a dataset $X=\{\bm{x}_{1},\bm{x}_{2},...\bm{x}_{N}\}$, a clustering algorithm aims to assign each sample $\bm{x}_{i}$ to a group such that a human-designed cost is minimized. The most frequently used clustering method for discrete speech tokens is k-means clustering [41], e.g. in GSLM [4]. K-means is a clustering algorithm based on Euclidean distances. Its training process iteratively assigns each data sample to the nearest centroid, and moves cluster centroids till convergence, with a pre-defined cluster number $k$. After training, the centroids form the codebook, and new data can be quantized to the index of the nearest centroid. In practice, centroids are usually initialized with the k-means++ algorithm [42] for better convergence.
Hierarchical agglomerative clustering has also been used in discrete speech tokens, which iteratively merges the closest clusters. It is usually applied after k-means to reduce the number of clusters [43, 44]. Other clustering algorithms are less explored in the context of discrete speech tokens, thus will not be covered here.
@@ II-B Vector Quantization
Although clustering methods are easy to implement and apply, they are often an isolate process, thus cannot be optimized together with other neural network modules. Instead, vector quantization (VQ), initially introduced by [45], enables a learnable network module that allows gradients to pass through when producing discrete representations. Autoencoders with a VQ module is termed VQ-VAE [46]. There are multiple ways a VQ module can be realized:
@@ II-B1 K-means VQ
Like k-means clustering, k-means VQ method finds the code-vector closest to the input in Euclidean space, i.e.
@@
As the $\arg\min$ operation is not differentiable, a straight-through estimator (STE) [47] is usually applied that grafts the gradient on $q(\bm{x})$ to $\bm{x}$ itself. Suppose some loss should numerically be calculated by $\mathcal{L}=\mathcal{L}(q(\bm{x}))$, then the STE performs gradient grafting by changing the input to $\mathcal{L}=\mathcal{L}(\bm{x}+\operatorname{sg}(q(\bm{x})-\bm{x}))$ where $\operatorname{sg(\cdot)}$ means the stop gradient operator. In this way, the loss function is still calculated by the value of $q(\bm{x})$, but gradients that should be placed on $q(\bm{x})$ is now grafted to $\bm{x}$ itself.
Auxiliary loss functions are often used together with k-means VQ: commitment loss $\mathcal{L}_{\text{cmt}}=\|\operatorname{sg}(\bm{c}_{k})-\bm{x}\|^{2}$ and codebook loss $\mathcal{L}_{\text{code}}=\|\operatorname{sg}(\bm{x})-\bm{c}_{k}\|^{2}$. The commitment loss pushes the continuous input $\bm{x}$ towards the closest codebook entry, while the codebook loss pushes the latter towards the former and updates the code-vector $\bm{c}_{k}$. The two loss terms are weighted by different factors to put different optimization strengths on $\bm{x}$ and $\bm{c}_{k}$, as pushing $\bm{c}_{k}$ towards $\bm{x}$ is an easier task. It is also common to replace $\mathcal{L}_{\text{code}}$ with exponential moving average (EMA) to update the codebook instead [48]. Although EMA does not rely on explicit loss functions, it achieves a similar goal that gradually merges the continuous input $\bm{x}$ into the code-vector $\bm{c}_{k}$.
VQ in the high-dimensional spaces is known to suffer from codebook collapse, a phenomenon where only a portion of codebook is active or the codebook usage is highly imbalanced [49, 50]. To improve the utilization of codebook entries, random replacement (as known as codebook expiration) is sometimes applied [50] on code-vectors that has remained inactive for a long time. Another direct solution is to add an additional auxiliary constraint, such as entropy penalty, to promote a more uniform distribution on the codebook [51, 52]. Architectural improvements have been proposed, like factorized codebook lookup in low-dimensional space [53], and adding a linear projection layer to simultaneously update all code-vectors [54].
@@ II-B2 Gumbel VQ
Instead of quantizing by Euclidean distance, another choice is probability. Gumbel VQ [55] uses Gumbel-Softmax as a proxy distribution for traditional Softmax to allow differentiable sampling. Given input $\bm{x}$ and a codebook of size $V$, a transform is applied on $\bm{x}$ into $V$ logits: $\bm{l}=h(\bm{x})\in\mathbb{R}^{V}$. In inference, quantization is then performed by choosing the index with the largest logit, i.e. $q(\bm{x})=\arg\max_{i}\left\{\bm{l}^{(i)}\right\}$. In training, samples are drawn from the categorical distribution implied by $\bm{l}$ for the subsequent neural networks. To achieve efficient sampling and let gradients pass through, Gumbel trick is used:
@@
where Eq.(2) samples Gumbel noise $\bm{v}$ element-wise, and Eq.(3) calculates Gumbel-Softmax distribution $\bm{s}$ with a temperature $\tau$. The forward pass simply use $j=\arg\max_{i}\{\bm{s}^{(i)}\}$ as the sampled index, but the true gradient of Gumbel-Softmax is used in backward pass. In other words, the gradient on the one-hot distribution corresponding to $j$ is grafted to $\bm{s}$ as an approximate. The temperature $\tau$ balances the approximation accuracy and difficulty of gradient passing: a lower $\tau$ results in sharper $\bm{s}$ and thus more accurate gradient estimate, but higher gradient variances [55]. In practice, $\tau$ is usually annealed from high to low [55, 30]. The transform $h(\cdot)$ is usually parameterized as neural networks, or negatively proportional to Euclidean distances [56].
After quantization, code-vector $\bm{c}_{i}$ with $i=q(\bm{x})$ is fed to subsequent networks. Gumbel VQ does not require additional losses for quantization, since code-vectors can directly be learned with gradients and do not need to be pushed towards $\bm{x}$ like k-means VQ.
@@ II-B3 Finite Scalar Quantization
As mentioned before, VQ methods based on code-vector assignment suffer from low codebook usage. Despite the introduction of many different mechanisms, this remains a crucial challenge. Finite scalar quantization (FSQ) [57] is an alternative to perform quantization in scalar domain. FSQ quantizes each dimension of a vector $\bm{x}$ into $L$ levels. For the $i$-th dimension $\bm{x}^{(i)}$, FSQ transforms the values to into limited range and then rounds to integers, i.e.
@@
The quantized values are thus integers ranging from $-\lfloor L/2\rfloor$ to $\lfloor L/2\rfloor$ [^2] Following [57], this is the symmetric case for $L$ being odd. When $L$ is even, there is an offset before rounding to obtain asymmetric quantized values.. For a $d$-dimensional vector $\bm{x}$, there are $L^{d}$ possible quantization outcomes. Common choices are $L\geq 5$ and $d\leq 10$, so FSQ usually has a much smaller hidden dimension than VQ (where usually $d\geq 100$). STE is applied to pass gradients. As quantization is simply done via rounding to integers, there is no explicit codebooks associated with the FSQ process. FSQ is reported to have better codebook usage [^3] Although there is no longer a codebook associated with code-vectors, codebook usage can still be measured among all possible $V=L^{d}$ outcomes. especially for a large $V$ compared to VQ methods, without the need for auxiliary losses.
@@ II-B4 Other VQ Tricks
In many cases, a single VQ module suffers from a highly-limited representation space, thus results in poor performance compared to continuous counterparts. There are two widely-used VQ tricks that introduce multiple VQ quantizers to expand the quantized representation space, as shown in Fig.3:
@@
    Residual VQ (RVQ), also known as multi-stage quantization [60]. It adopts a serial approach that iteratively quantizes the residual of the last quantizer. Similar to GVQ, RVQ also has multiple quantizers. For the $i$-th quantizer $q_{i}$, denote its input as $\bm{x}_{i}$ and the output code-vector as $\bm{c}_{q_{i}(\bm{x}_{i})}$, then the residual is defined as $\bm{x}_{i+1}=\bm{x}_{i}-\bm{c}_{q_{i}(\bm{x}_{i})}$, i.e. the “quantization margins”. The quantized results from all $q_{i}$ are finally summed as an approximation of original $\bm{x}$. In this way, information in the codebooks is supposed to follow a coarse-to-fine order, and more details in the original $\bm{x}$ can be preserved. It is used in various speech codecs [61, 62, 18], for instance.
@@
Note that RVQ naturally produces an order of importance in residual layers, while all quantizers in GVQ are equally important. There is also a trick called “nested dropout” [65] that assigns an importance order to GVQ. This works by manually define an order of quantizers and randomly dropping-out the last few quantizers in training.
@@ III Speech Tokenization Methods: Acoustic Tokens
We summarize the quantization method, number of quantizers $Q$, frame rate $F$, vocabulary size $V$ for each quantizer, and the resulting bitrate of famous neural acoustic speech tokens in Table. I.

### v2

@@ I Introduction
The rapid advancement of large language models (LLMs) in natural language processing has revolutionized speech generation tasks [1, 2], with speech being tokenized and modeled using decoder-only Transformers [3]. Efforts starting from GSLM [4] and AudioLM [5] aim to develop text-free spoken LLMs, akin to how current LLM-powered chatbots enable text-based interactions. Other works, including VALL-E [6] and VioLA [7], extend this approach to conditional speech generation tasks, such as zero-shot text-to-speech and speech translation. However, this paradigm requires data to be tokenized, as LLMs typically process discrete data only. Textual tokens naturally meet this requirement because they are designed as discrete units separated by clear boundaries, whereas raw speech signals are continuous and boundary-less. Therefore, a necessary step before applying speech data to LLM is the tokenization of speech, whose goal is:
To convert long speech waveforms into short discrete tokens for downstream tasks. These tokens should be compatible with the underlying textual representations, especially for language modeling approaches targeted at speech.
@@
As a result, significant efforts have been directed towards developing efficient and powerful speech tokenization methods. Generally, these methods are based on two distinct principles, giving rise to two types of speech tokens: acoustic tokens and semantic tokens. Acoustic tokens are derived from neural codecs designed to encode speech at a low bitrate while preserving as much information as possible. In contrast, semantic tokens originate from speech self-supervised learning (SSL) [8], which aims to learn a more phonetic or semantic representation space, making it easier for speech recognition. These two nearly independent lines of research magically intersect in the context of language modeling for speech. Now, there are also efforts that try to design a speech tokenizer that accomplishes the two objectives simultaneously [9, 10]. Consequently, speech tokenization has become a core problem of speech processing under the new paradigm, with versatile downstream applications, as shown in Fig.1.
Despite the rapid development and numerous recent works, a comprehensive taxonomy of methodologies in discrete speech tokens has not been clearly constructed. Existing reviews [8, 11, 12, 1, 13, 2] in this field overlook the diverse categories and methodologies in both acoustic and semantic tokens. For example, [1, 2] focus primarily on methods in spoken language modeling, providing only brief descriptions of some speech tokens used in existing models. The taxonomy of neural audio codecs has been summarized in [12, 14], but the realm of semantic tokens is still overlooked. In this review, we provide a comprehensive overview of the concepts, methods, and characteristics of various types of discrete speech tokens, with their applications in spoken language understanding, speech generation, and spoken dialogue models. We hope that through this review, the community can have a clear understanding of the current development and key technologies of discrete speech tokens, so as to promote further research in the future.
@@
  • This review is the first to focus specifically on discrete speech tokens with sufficient depth in the LLM era.
@@
The structure of this review is shown in Fig.2. Following [5, 15, 16], we classify discrete speech tokens into acoustic and semantic tokens based on their principles. We will characterize the two types of tokens both by conceptual descriptions and unified experimental comparisons.
@@ II Pre-requisites: Discrete Representation Learning
Discrete speech tokens are obtained through the quantization of continuous representations, which is usually achieved by offline clustering or online vector quantization algorithms. This section provides a concise overview of the existing quantization methods commonly used in discrete speech tokens.
Denote $\bm{x}\in\mathbb{R}^{d}$ as a vector in the $d$-dimensional continuous space. A quantization process $q$ transforms $\bm{x}$ into a discrete token in a finite set, i.e. $q(\bm{x}):\mathbb{R}^{d}\to\{1,2,...,V\}$ where $V$ is the vocabulary size. The output tokens are sometimes referred to as indexes in the finite $V$-cardinal set. The function $q$ is usually associated with a codebook $\mathcal{C}=\{\bm{c}_{1},\bm{c}_{2},...,\bm{c}_{V}\}$ where every code-vector $\bm{c}_{i}\in\mathbb{R}^{d}$ corresponds to the $i$-th token. The code-vectors are representations of tokens in the original $d$-dimensional space. As $V$ elements can be encoded using $\lceil\log_{2}V\rceil$ raw bits [^1] We use $\lceil z\rceil$ to denote the ceiling of a scalar $z$, i.e., the smallest integer greater than or equal to $z$. Similarly, $\lfloor z\rfloor$ denotes the floor of $z$, i.e., the largest integer less than or equal to $z$., quantization often compresses the cost for data storage and transmission to a great extent.
@@ II-A Offline Clustering
Clustering is a simple approach for quantization. Given a dataset $X=\{\bm{x}_{1},\bm{x}_{2},...\bm{x}_{N}\}$, a clustering algorithm aims to assign each sample $\bm{x}_{i}$ to a group such that some cost is minimized. The most frequently used clustering method for discrete speech tokens is k-means clustering [42], e.g. in GSLM [4]. K-means is a clustering algorithm based on Euclidean distances. Its training process iteratively assigns each data sample to the nearest centroid, and moves cluster centroids till convergence, with a pre-defined number of clusters. After training, the centroids form the codebook, and new data can be quantized to the index of the nearest centroid in this Voronoi partition. In practice, centroids are usually initialized with the k-means++ algorithm [43] for better convergence.
Hierarchical agglomerative clustering has also been used in discrete speech tokens, which iteratively merges the closest clusters. It is usually applied after k-means to reduce the number of clusters [44, 45]. Other clustering algorithms are less explored in the context of discrete speech tokens.
@@ II-B Vector Quantization
Clustering is often an isolate process, thus cannot be optimized together with other neural network modules. Instead, vector quantization (VQ) [46] enables a learnable network module that allows gradients to pass through when producing discrete representations. Autoencoders with a VQ module is termed VQ-VAE [47]. There are multiple VQ methods:
@@ II-B1 K-means VQ
Like k-means clustering, k-means VQ method finds the code-vector closest to the input, i.e.
@@
Then, code-vector $\bm{c}_{k}\triangleq\bm{c}_{q(\mathbf{x})}$ is fed to subsequent networks. As the $\min$ operation is not differentiable, straight-through estimators (STEs) [48] are usually applied to graft gradients, i.e. $\operatorname{STE}(\bm{c}_{k},\bm{x})=\bm{x}+\operatorname{sg}(\bm{c}_{k}-\bm{x})$ where $\operatorname{sg(\cdot)}$ stops tracking gradients. In this way, the input value to subsequent networks is still $\bm{c}_{k}$, but gradients are grafted to $\bm{x}$ in back propagation.
Auxiliary loss functions are often used together with k-means VQ [47]: commitment loss $\mathcal{L}_{\text{cmt}}=\|\operatorname{sg}(\bm{c}_{k})-\bm{x}\|^{2}$ and codebook loss $\mathcal{L}_{\text{code}}=\|\operatorname{sg}(\bm{x})-\bm{c}_{k}\|^{2}$. The commitment loss pushes the continuous input $\bm{x}$ towards the closest codebook entry, while the codebook loss does the opposite and updates the code-vector $\bm{c}_{k}$. The two loss terms are weighted by different factors to put different optimization strengths on $\bm{x}$ and $\bm{c}_{k}$, as pushing $\bm{c}_{k}$ towards $\bm{x}$ is an easier task. It is also common to replace $\mathcal{L}_{\text{code}}$ with exponential moving average (EMA) to update the codebook instead [49], which does not rely on explicit loss functions.
VQ in high-dimensional spaces is known to suffer from codebook collapse, where the codebook usage is highly imbalanced [50, 51]. To improve the utilization of codebook, random replacement (as known as codebook expiration) can be applied [51] on code-vectors that have remained inactive for a long time. Other solutions include additional auxiliary constraints such as entropy penalty [52, 53], factorized codebook lookup in low-dimensional space [54], and adding a linear projection to update all code-vectors together [55].
@@ II-B2 Gumbel VQ
Instead of quantizing by Euclidean distance, another choice is by probability. Gumbel VQ [56] uses Gumbel-Softmax as a proxy distribution for traditional Softmax to allow differentiable sampling. Given input $\bm{x}$ and a codebook of size $V$, a transform $h(\cdot)$ is applied on $\bm{x}$ into $V$ logits: $\bm{l}=h(\bm{x})\in\mathbb{R}^{V}$. In inference, quantization is performed by choosing the index with the largest logit, i.e. $q(\bm{x})=\arg\max_{i}\left\{\bm{l}^{(i)}\right\}$. In training, samples are drawn from the categorical distribution implied by $\bm{l}$ for the subsequent neural networks. To achieve efficient sampling and let gradients pass through, Gumbel trick is used:
@@
where Eq.(2) samples Gumbel noise $\bm{v}$ element-wise, and Eq.(3) calculates Gumbel-Softmax distribution $\bm{s}$ with a temperature $\tau$. The forward pass simply use $j=\arg\max_{i}\{\bm{s}^{(i)}\}$ as the sampled index, but the true gradient of Gumbel-Softmax is used in backward pass. In other words, the gradient on the one-hot distribution corresponding to $j$ is grafted to $\bm{s}$ as an approximate. The temperature $\tau$ balances the approximation accuracy and gradient variances. The transform $h(\cdot)$ is usually parameterized as neural networks, or negatively proportional to Euclidean distances [57].
After quantization, code-vector $\bm{c}_{k}$ with $k=q(\bm{x})$ is fed to subsequent networks. Gumbel VQ does not require additional losses, since code-vectors can be directly learned with gradients and do not need to be pushed towards $\bm{x}$.
@@ II-B3 Finite Scalar Quantization
As mentioned before, VQ methods based on code-vector assignment usually suffer from codebook collapse. Despite many efforts, this remains a crucial challenge. Finite scalar quantization (FSQ) [58] is an alternative to perform quantization in scalar domain. FSQ quantizes each dimension of a vector $\bm{x}$ into $L$ levels. For the $i$-th dimension $\bm{x}^{(i)}$, FSQ transforms the values to into limited range and then rounds to integers, i.e.
@@
The quantized values are thus integers ranging from $-\lfloor L/2\rfloor$ to $\lfloor L/2\rfloor$ [^2] Following [58], this is the symmetric case for $L$ being odd. When $L$ is even, there is an offset before rounding to obtain asymmetric quantized values.. For a $d$-dimensional vector $\bm{x}$, there are $L^{d}$ possible quantization outcomes. Hence, FSQ usually requires a much smaller hidden dimension than VQ. STE is applied to pass gradients. As quantization is simply done via rounding to integers, there is no explicit codebooks associated with the FSQ process. FSQ is reported to have better codebook usage [^3] Although there is no longer a codebook associated with code-vectors, codebook usage can still be measured among all possible $V=L^{d}$ outcomes. especially for a large $V$ compared to VQ methods, without the need for auxiliary losses.
@@ II-B4 Other VQ Tricks
In many cases, a single VQ module suffers from a highly-limited representation space, thus results in poor performance compared to continuous counterparts. There are some widely-used VQ tricks that introduce multiple quantizers to refine the quantized space, as shown in Fig.3:
@@
    Residual VQ (RVQ), also known as multi-stage quantization [61]. It adopts a serial approach that iteratively quantizes the residual of the last quantizer. Similar to GVQ, RVQ also has multiple quantizers. For the $i$-th quantizer $q_{i}$ with input $\bm{x}_{i}$ and output code-vector $\bm{c}_{k}$, the residual is defined as $\bm{x}_{i+1}=\bm{x}_{i}-\bm{c}_{k}$. The outputs from all $q_{i}$ are finally summed as the quantized result of $\bm{x}$. In this way, information in the codebooks is supposed to follow a coarse-to-fine order, and more details in the original $\bm{x}$ can be preserved than a plain quantizer. It is used in various speech codecs [62, 63, 19], for instance.
@@
Note that RVQ naturally produces an order of importance in residual layers, while all quantizers in GVQ are equally important. Such order of importance can also be enforced in GVQ by a “nested dropout” trick [66].
@@ III Speech Tokenization Methods: Acoustic Tokens
We list the VQ method, number of quantizers $Q$, frame rate $F$, vocabulary size $V$ for each quantizer, and the resulting bitrate of existing neural acoustic speech tokens in Table.I.